Part A It takes 55.0 J to raise the temperature of an 11.5 g piece of unknown metal from 13.0° Cto 244 °C. What is the specific heat for the metal? Express your answer with the appropriate units. View Available Hint(s) 因国1tỉ l ● x-10n @白 Units G-1 Value Submit
media%2F580%2F58078b0f-910f-4f1f-a637-c7
media%2Fbf0%2Fbf019e6d-3279-423c-b760-2b
Answer #1

Part A

Specific heat of metal (Cs) = q/m*DT

                            = 55.0/(11.5*(24.4-13.0))

                            = 0.42 j/g.c
Part B

molar heat capacity of silver(Cm) = 25.35 j/mol.k

molar heat capacity of silver(Cm) = q/n*DT

n = no of mol of silver = 11.5/108 = 0.1065 mol

DT = 10.1 C

q = heat energy = ?

25.35 = q/(0.1065*10.1)

q = 27.27 j

Part C

specific heat of silver(Cs) = Cm/M

Cm = molar heat capacity of silver(Cm) = 25.35 j/mol.k

M = molarmass of silver = 108 g/mol

Cs= 25.35/108

   = 0.235 j/g.mol